### v1

@@ Abstract
We find that the spontaneous and collective emissions have a large influence on the excitation of two-level absorbers (atoms, molecules) interacting in resonance with the plasmonic mode near the metal nanoparticle. The spontaneous and collective emissions restrict the absorption enhancement by plasmonic mode and make the enhancement possible only with a fast, on a pico-second time scale, population relaxation of the upper absorbing states. Conditions for the maximum of the plasmon-enhanced absorption in the presence of spontaneous and collective emissions are found. We show that the nanoparticle-absorber interaction and the collective emission cause the bistability in the plasmon-enhanced absorption.
I Introduction
@@
The localized plasmon resonances (LPR) excitation leads to the electric field enhancement near metal nanoparticles Maier (2007); Pelton, Aizpurua, and Bryant (2008); Amendola et al. (2017); Tanabe (2008). Such a "plasmonic" field enhancement finds various applications in optical sensors Li, Cushing, and Wu (2015), bio-medicine Liao, Nehl, and Hafner (2006), photovoltaics Jang et al. (2016), photo-catalysis Zhang et al. (2013) and other areas of physics and chemistry Maier (2007); Barbillon (2019). A new branch of chemistry - plasmonic chemistry was developed concerning the chemical reactions Zhan et al. (2023); Cortés (2017) stimulated by the electric field enhanced near plasmonic nanoparticles.
@@
A photochemical reaction is initiated by the absorption of light and the creation of transient molecules in excited states whose properties differ greatly from the original molecules. These new chemical species can fall apart, change to new structures, combine other molecules, or transfer electrons, hydrogen atoms, protons, or their electronic excitation energy to other molecules Fleming, Krueger, and Longworth (2023). So the optimization of the plasmonic photo-chemical reactions is important with the searching for conditions of the maximum field and absorption enhancement near metal nanoparticles Kelly et al. (2003). This paper contributes to such an optimization.
A large electric field enhancement is well-known near the sharp features of nanostructures, as "hot spots" Díaz-Núñez et al. (2019), or in a small gap between nanoparticles Kim et al. (2010). However, it may be difficult to make a sufficiently large number of "hot spots" or small gaps in nanostructures. So in this paper, we do not consider the field enhancement by "hot spots" or in gaps between nanoparticles. We restrict ourselves by the absorption enhancement in the near field of metal nanoparticles with the excited LPR.
@@
We will pay attention on the spontaneous and collective emissions, which significantly affect the resonant absorption near plasmonic nanoparticles and cause new physical effects. Spontaneous emission near metal nanoparticles has been investigated, for example, in Blanco and García de Abajo (2004). Collective effects as a super-radiance have been studied for the resonant absorbers, as atoms or molecules, in the near zone of the plasmonic particle Shahbazyan and Pustovit (2011); Protsenko and Uskov (2015); Protsenko et al. (2017). However, the absorption enhancement near the nanoparticle with an account for the spontaneous and collective emissions deserves more studies. The influence of spontaneous emission and collective effects on absorption is important for plasmonic chemistry, particularly, in the relation with recently studied plasmon-induced direct resonance energy transfer (PIRET) processes You, Ramakrishna, and Seideman (2018). A part of studies of the plasmon-enhanced absorption with the spontaneous and collective emissions to the LPR mode, with the direct energy transfer from the plasminic nanoparticle to the absorbing molecule, is done in the present paper.
@@
Here we use the model for the resonant absorption near plasmonic nanoparticles similar to the laser model. The plasmonic nanoparticle makes a cavity for the electric field similar to the laser cavity, as in the plasmonic nanolaser Protsenko et al. (2005, 2008); Protsenko (2012). A part of our approach is similar to the analysis of the spontaneous emission and collective effects in lasers André et al. (2019); Protsenko and Uskov (2022).
@@
The plasmonic mode provides good conditions for collective effects in the resonant absorption. Collective effects in the resonant medium are significant in the low-quality cavity André et al. (2019); Protsenko and Uskov (2022); Piccione, Felicetti, and Bellomo (2022). The LPR mode has a low quality, appropriate for collective effects. The particle surface is a good place for the resonant absorbers (atoms, molecules, q-dots, etc.).
@@
In this paper, we show that the spontaneous emission into the LPR mode and collective interaction of absorbers through the plasmonic electric field significantly affect the absorber excitation in the vicinity of metal nanoparticles and lead to bistability in the excitation. We will find conditions for the best absorber excitation and bistability. These results are important for all branches of plasmonic chemistry, photovoltaic, and all applications related to the resonant excitation of absorbers near plasmonic nanoparticles.
@@
We derive basic equations quantum-mechanically as Heisenberg operator equations. In the present paper, we solve these equations in the semi-classical approximation as it is common, for example, in the laser theory Sargent, Scully, and Lamb (1974). Fully quantum analysis of the equations will be done in the future.
@@
We carry out calculations analytically. The analytical approach lets to find expressions illustrating the physical processes at the absorber excitation and the influence of the plasmonic particle on important parameters such as the polarization and the population of states of absorbers.
We consider the increase of the absorber excited state population as a criterion of absorption enhancement by the LPR.
In the next section, we describe the model, Hamiltonian, and equations of motion.
@@
The expressions, characterizing the stationary absorption, as the absorption coefficients and the energy balance equation, are presented in section III. Several physical processes at the resonant absorption near the metal nanoparticle are explained with the help of equations found in section III.
@@
Section IV shows conditions for the best plasmon-enhanced absorption and describes the bistability.
@@
Examples of the absorption enhancement and the bistability are presented in section V at typical values of the nanoparticle and the absorber parameters.
Results are discussed in section VI, and the paper is finalized in the conclusion section VII.
II The model, Hamiltonian and equations of motion
II.1 The model
@@
Suppose, a metal nanoparticle interacts with $i=1...N_{m}$ identical absorbers, as molecules or atoms, being on the distance $r$ from the particle center on the dielectric shell as shown in Fig. 1. The absorber transition frequency is resonant with the dipole LPR mode of the nanoparticle.
@@
Figure 1: Coordinate system, the nanoparticle 1 and the resonant absorber $i$ (a molecule or an atom) in the external electric field $\hat{\vec{E}}$. $r$ is the distance, and $\vec{e}_{ri}$ is the unit vector from the nanoparticle center to the absorber. $\left|00\right>_{i}$ is the absorber ground state. The dipole momentum with the matrix element $d$ for the absorber transition from the excited state $\left|01\right>_{i}$ (or $\left|10\right>_{i}$) to $\left|00\right>_{i}$ is parallel to the unit vector $\vec{e}$ of the external field polarization (or to the vector $\vec{e}_{i}$). The absorber state with the transition dipole momentum parallel to the axis $OO^{\prime}$ is not excited; such transition does not interact with the external and the nanoparticle fields. The insert shows the resonant absorbers randomly distributed on the nanoparticle shell.
Such resonance can be achieved for the nanoparticle of appropriate shape Kelly et al. (2003). For simplicity, we consider a spherical nanoparticle.
@@
The particle and absorbers interact with the external monochromatic electric field of the amplitude $\hat{\vec{E}}$ linearly polarized along the unit vector $\vec{e}$. The field $\hat{\vec{E}}$ is resonant with the LPR and the absorber transitions.
@@
We consider the $i$-th absorber in the Cartesian coordinate system with the origin in the particle center. One coordinate axis is parallel to $\vec{e}$, another axis with the unit vector $\vec{e}_{i}$ is in the plane made by $\vec{e}$ and the unit vector $\vec{e}_{ri}$ directed from the nanoparticle center to the i-th absorber. The angle $\theta_{i}$ is between $\vec{e}$ and $\vec{e}_{i}$, and the third coordinate axis is along the line $OO^{\prime}$. Two excited ${{\left\langle 10\right|}_{i}}$, ${{\left\langle 01\right|}_{i}}$ and the ground ${{\left\langle 00\right|}_{i}}$ states of i-th absorber participate in the interaction with electric fields near the nanoparticle. Such absorber states are, for example, the lowest unoccupied molecular orbital (LUMO) and the highest occupied molecular orbital (HOMO) Fukui, Yonezawa, and Shingu (2004).
@@
We will calculate the population of excited states of absorbers and examine how the interaction between the particle and absorbers affects the population.
@@
We suppose a large number $N_{m}\gg 1$ of absorbers randomly distributed on the nanoparticle shell as shown on the insert in Fig. 1. In the case of $N_{m}\gg 1$, the absorber-absorber interaction through the near field leads to only the self-broadening of the absorber’s transitions Royer (1975). We include the self-broadening in the absorber transition width $\gamma$ and do not describe the absorber-absorber interaction explicitly.
II.2 Hamiltonian
@@
Following Protsenko et al. (2005) we describe the nanoparticle as a dipole making harmonic oscillations in the resonant external electric field. We introduce the nanoparticle dipole momentum operator
@@
where $p$ is the dipole momentum matrix element, $\hat{b}$ is a Bose-operator describing collective (plasmonic) oscillations of the nanoparticle electrons in the electric field. Later we express $p$ through the nanoparticle polarizability Protsenko et al. (2005). For symmetry reasons, we suppose the same directions for the nanoparticle dipole and the external field polarization.
The dipole momentum operator of the $i$-th absorber resonant transitions is
${{\hat{\vec{d}}}_{i}}=d\left[\left({{{\hat{\sigma}}}_{i}}\vec{e}+{{{\hat{\sigma}}}_{\bot i}}{{{\vec{e}}}_{i}}\right){{e}^{-i{{\omega}_{0}}t}}+h.c.\right].$ (2)
Here $d$ is the dipole transition matrix element; absorber operators in Dirac notations ${{\hat{\sigma}}_{i}}={{\left|00\right\rangle}_{i}}{{\left\langle 01\right|}_{i}}$, (or ${{{\hat{\sigma}}}_{\bot i}}={{\left|00\right\rangle}_{i}}{{\left\langle 10\right|}_{i}}$) are transition operators from the high $\left|01\right>_{i}$ (or $\left|10\right>_{i}$) to the low $\left|00\right>_{i}$ states of the absorber $i$ parallel to $\vec{e}$, (or $\vec{e}_{i}$): ${{{\hat{\sigma}}}_{i}}\left|01\right>_{i}=\left|00\right>_{i}$ and ${{{\hat{\sigma}}}_{\perp i}}\left|10\right>_{i}=\left|00\right>_{i}$.
Hamiltonian of the particle and $N_{m}$ absorbers in the electric field $\hat{\vec{E}}$ is
@@
The first term in (3) describes free oscillations of dipoles; the second term is for the dipole interaction of the particle and absorbers with the external field; the third term ${{\hat{V}}_{int}}$ describes the interaction of the particle and absorbers through the nanoparticle near field (the LPR field) and the last term in (3) describes the dissipation and the interaction of the particle and absorbers with baths.
We suppose that the distance $r$ between absorbers and the particle center is much smaller than the LPR wavelength. So the particle-absorber interaction ${{\hat{V}}_{int}}$ is a quasi-static dipole-dipole interaction. We take the dipole-dipole quasi-static interaction energy from Landau and Lifshitz (1982), replace classical dipoles with operators, and write
@@
where ${{{\vec{e}}}_{ri}}$ is the unit vector along the direction from the nanoparticle center to i-th absorber as shown in Fig. 1.
@@
where $\hat{a}$ is the field amplitude Bose-operator, ${{E}_{0}}=\eta^{-1}\sqrt{{2\pi\hbar{{\omega}_{0}}}/{V}}$, $V$ is the field quantization volume ($V$ does not come to final results Lifshitz, Pitaevskii, and Berestetskii (1982)), $\eta$ is the refractive index of the nanoparticle environment Tkalya (2000). We suppose that distances between absorbers and the particle center are much smaller than the external field wavelength, so the phase (taken zero) of the external field is the same for all absorbers and the particle.
We substitute expressions (1), (2) into the sum in Eq (4), drop the fast oscillating terms $\sim e^{\pm 2i\omega_{0}t}$ and represent
@@
Expression (6) can be simplified at a large number of absorbers $N_{m}\gg 1$. Summing over $i$ we approximate ${{\cos}^{2}}{{\theta}_{i}}\approx 1/2$ and neglect the product ${{{\hat{\sigma}}}_{\bot i}}\cos{{\theta}_{i}}\sin{{\theta}_{i}}$ repeatedly changing the sign with $i$. After such a simplification of Eq. (6) the Hamiltonian (3) written in the interaction picture Landau and Lifshitz (2013) and RWA approximation Sargent, Scully, and Lamb (1974) reads
@@
where ${{\Omega}_{p}}={p{{E}_{0}}}/{\hbar}$, ${{\Omega}_{m}}={d{{E}_{0}}}/{\hbar}$ and ${{\Omega}_{int}}={pd}/2\hbar{{r}^{3}}$. The Hamiltonian (7) corresponds to the approximation when the polarization of absorber’s transitions is parallel to the external field polarization.
II.3 Equations of motion
We write Heisenberg equations of motion $i\hbar\dot{\hat{A}}=\left[\hat{A},H\right]$ where $\hat{A}$ is an operator from Hamiltonian (7). The equation for the plasmonic oscillation Bose-operator $\hat{b}$ is

### v2

@@ Abstract
We find that the spontaneous and collective emissions have a strong influence on the excitation of two-level absorbers (atoms, molecules) interacting in resonance with the plasmonic mode near the metal nanoparticle. The spontaneous and collective emissions limit the absorption enhancement by the plasmonic mode and make the enhancement possible only with a fast, picosecond population relaxation of the upper absorbing states. Conditions for the maximum of plasmon-enhanced absorption in the presence of spontaneous and collective emissions are found. The nonlinearity in the nanoparticle-absorber interaction and in collective emission causes the bistability in the plasmon-enhanced absorption at high external field intensities and the plasmonic mode excitation.
keywords:
plasmonics , nanoparticle , absorption
PACS:
78.67.Bf , 73.20.Mf , 78.67.Qa , 73.22.Dj
journal: Photonics and nanostructures …
[…]
[inst1]organization=P.N.Lebedev Physical Institute of the RAS,addressline=Leninsky prospect 53, city=Moscow, postcode=119991, country=Russia
[…]
[inst2]organization=ITMO University,addressline=Kronverksky Pr. 49, city=St. Petersburg, postcode=197101, country=Russia
1 Introduction
@@
Excitation of localised plasmon resonances (LPR) leads to electric field enhancement in the vicinity of metal nanoparticles [1, 2, 3, 4]. Such a ”plasmonic” field enhancement finds various applications in optical sensors [5], bio-medicine [6], photovoltaics [7], photocatalysis [8] and other areas of physics and chemistry [1, 9]. A new branch of chemistry - plasmonic chemistry - has been developed around the chemical reactions [10, 11] stimulated by the electric field enhancement near plasmonic nanoparticles.
@@
A photochemical reaction is initiated by the absorption of light and the creation of transient molecules in excited states whose properties differ greatly from the original molecules. These new chemical species can fall apart, change into new structures, combine other molecules, or transfer electrons, hydrogen atoms, protons, or their electronic excitation energy to other molecules [12]. Therefore, it is important to optimise the plasmonic photochemical reactions by searching for conditions of maximum field and absorption enhancement near metal nanoparticles [13]. This work contributes to such optimization.
A large electric field enhancement is known to occur near the sharp features of nanostructures, known as ’hot spots’ [14], or in a small gap between nanoparticles [15]. A plasmonic gap field mode has been shown to have a high potential for achieving ultra-strong light absorption, the magnitude of which can be relatively easily controlled [16, 17, 18, 19]. In this paper we consider, as a first approximation, a much simpler system, limited to the near-field absorption enhancement of the isolated metal nanoparticle with the excited LPR. The approach and results of the present simple model will help to analyse the perspective but more complicated problem of absorption enhancement in plasmonic gaps.
@@
We will focus on the spontaneous and collective emissions, which significantly affect the resonant absorption near plasmonic nanoparticles and lead to new physical effects. Spontaneous emission near metal nanoparticles has been studied in many papers, for example in [20]. Collective effects in the form of superradiance have been studied for the resonant absorbers, such as atoms or molecules, in the vicinity of the plasmonic particle [21, 22, 23]. However, the absorption enhancement in the vicinity of the nanoparticle, taking into account the spontaneous and collective emissions, deserves further study. The influence of spontaneous emission and collective effects on absorption is important for plasmonic chemistry, especially in the context of the recently studied plasmon-induced direct resonance energy transfer (PIRET) processes [24]. Part of the study of plasmon-enhanced absorption with spontaneous and collective emission in the LPR mode, with direct energy transfer from the plasmonic nanoparticle to the absorbing molecule, is carried out in the present work.
@@
Here we use the model for resonant absorption near plasmonic nanoparticles, similar to the laser model. The plasmonic nanoparticle forms a cavity for the electric field similar to the laser cavity as in the plasmonic nanolaser [25, 26, 27]. Part of our approach is similar to the analysis of spontaneous emission and collective effects in lasers [28, 29].
@@
The plasmonic mode provides good conditions for collective effects in the resonant absorption. Collective effects in the resonant medium are significant in the low quality cavity [28, 29, 30]. The LPR mode is of low quality. The particle surface is a good location for the resonant absorbers (atoms, molecules, q-dots, etc.).
@@
In this paper, we show that the spontaneous emission in the LPR mode and the collective interaction of absorbers through the plasmonic electric field significantly affect the absorber excitation in the vicinity of metal nanoparticles and lead to bistability in the excitation. We will find conditions for absorber excitation enhancement and bistability. These results are important for all branches of plasmonic chemistry, photovoltaics and all applications related to resonant excitation of absorbers near plasmonic nanoparticles.
@@
We derive the fundamental equations quantum mechanically as Heisenberg operator equations. In the present work, we solve these equations in the semi-classical approximation, as it is used for example in laser theory [31]. A full quantum analysis of the equations will be performed in the future.
@@
We perform analytical calculations. The analytical approach allows us to find expressions that illustrate the physical processes at the excitation of the absorber and the influence of the plasmonic particle on important parameters such as the polarisation and the population of states of the absorber.
We consider the increase in the absorber excited state population as a criterion for absorption enhancement by the LPR. We describe the interaction of the dipole absorbers (molecules) with the dipole mode of the spherical plasmonic nanoparticle. According to Mie theory [32], the external field of the nanoparticle dipole mode is the field of the dipole in the nanoparticle center. Thus, the nanoparticle-absorber interaction is the dipole-dipole interaction as long as the high-order plasmonic mode excitation is neglected. We assume the resonance between the absorber transition and the frequency of the plasmonic dipole mode. The resonant dipole-dipole interaction at a short absorber-nanoparticle distance causes the well-known Förster nonradiative resonance energy transfer (FRET) [33]. Thus, in our work, the effect of collective and single absorber spontaneous emission in the plasmonic mode on the FRET process is highlighted.
We assume that the molecules are located on the nanoparticle dielectric shell with a thickness of at least 5 nm, which is sufficient to avoid strong absorber fluorescence quenching [34].
We assume that the nanoparticle radius is $13$ nm, so the total distance of $18$ nm between the nanoparticle center and the absorbers on the shell is much smaller than the optical wavelength of the resonant field $\sim 1~{}\mu$m, see the scheme in Fig. 1. We therefore use the quasi-static approximation for the absorber-nanoparticle dipole-dipole interaction.
We take into account the nonradiative and radiative [35] losses of the nanoparticle, including the losses associated with the collisions of the metal electrons with the nanoparticle surface [35, 36]. Thus, such a model based on the dipole-dipole interaction of the absorbers with the nanoparticle is a good approximation for our purposes. We note that more sophisticated models, such as macroscopic QED for quantum nanophotonics [37] and other approaches as discussed in the review [38], can be used for advanced investigation of the spontaneous and collective emission in plasmonic nanostructures outlined in this paper. A relatively simple model given below is useful for finding similarities in the collective emission phenomena in small lasers, LEDs and resonant absorber plasmonic structures.
In the next section, we describe the model, the Hamiltonian, and the equations of motion.
@@
The expressions characterising the steady-state absorption, such as the absorption coefficients and the energy balance equation, are presented in section 3. Several physical processes in the resonant absorption near the metal nanoparticle are explained with the help of the equations found in section 3.
@@
Section 4 shows the conditions for the best plasmon-enhanced absorption and describes the bistability.
@@
Examples of the absorption enhancement and the bistability are shown in section 5 for typical values of the nanoparticle and the absorber parameters.
The results are discussed in section 6 and the paper is concluded in section 7.
2 The model, Hamiltonian and equations of motion
2.1 The model
@@
Suppose, a metal nanoparticle interacts with $i=1...N_{m}$ identical absorbers, as molecules or atoms, located at the distance $r$ from the particle center on the dielectric shell as shown in Fig. 1. The transition frequency of the absorber is in resonance with the dipole LPR mode of the nanoparticle.
@@
Figure 1: Coordinate system, the nanoparticle 1 and the resonant absorber $i$ (a molecule or an atom) in the external electric field $\hat{\vec{E}}$. $r$ is the distance, and $\vec{e}_{ri}$ is the unit vector from the centre of the nanoparticle to the absorber. $\left|00\right>_{i}$ is the ground state of the absorber. The dipole momentum with the matrix element $d$ for the absorber transition from the excited state $\left|01\right>_{i}$ (or $\left|10\right>_{i}$) to $\left|00\right>_{i}$ is parallel to the unit vector $\vec{e}$ of the external field polarisation (or to the vector $\vec{e}_{i}$). The absorber state with the transition dipole momentum parallel to the axis $OO^{\prime}$ is not excited; such transition does not interact with the external and the nanoparticle fields. The inset shows the resonant absorbers randomly distributed on the nanoparticle shell.
Such resonance can be achieved for the nanoparticle of appropriate shape [13]. For simplicity, we will consider a spherical nanoparticle. A metallic nanoparticle with resonant absorbers/emitters on the dielectric shell is a well-known object in plasmonic theory and experiments. It has been considered and realised, for example, in the plasmonic nano-lasers [39, 40, 27], for the studies of the plasmonic superradiance [41, 22] and, most recently, in the discovery of the role of electron spillout in strong coupling of the plasmonic nanoparticle with the excitonic shell [42].
@@
The particle and the absorbers interact with the external monochromatic electric field of the amplitude $\hat{\vec{E}}$ linearly polarised along the unit vector $\vec{e}$. The field $\hat{\vec{E}}$ is in resonance with the LPR and the absorber transitions.
@@
We consider the $i$-th absorber in the Cartesian coordinate system with the origin at the particle centre. One coordinate axis is parallel to $\vec{e}$, another axis with the unit vector $\vec{e}_{i}$ is in the plane formed by $\vec{e}$ and the unit vector $\vec{e}_{ri}$ points from the nanoparticle centre to the i-th absorber. The angle $\theta_{i}$ is between $\vec{e}$ and $\vec{e}_{i}$, and the third axis of coordinate is along the line $OO^{\prime}$. Two excited ${{\left\langle 10\right|}_{i}}$, ${{\left\langle 01\right|}_{i}}$ and the ground ${{\left\langle 00\right|}_{i}}$ states of the i-th absorber participate interact with electric fields near the nanoparticle. Such absorber states are, for example, the lowest unoccupied molecular orbital (LUMO) and the highest occupied molecular orbital (HOMO) [43].
@@
We will calculate the population of excited states of the absorbers and study how the interaction between the particle and the absorbers affects the population.
@@
We assume a large number $N_{m}\gg 1$ of absorbers randomly distributed on the nanoparticle shell, as shown in the inset in Fig. 1. In the case of $N_{m}\gg 1$, the absorber-absorber interaction through the near-field leads only to the self-broadening of the absorber’s transitions [44]. We include the self-broadening in the absorber transition width $\gamma$ and do not explicitly describe the absorber-absorber interaction.
2.2 Hamiltonian
@@
Following [25], we describe the nanoparticle as a dipole that undergoes harmonic oscillations in the resonant external electric field. We introduce the nanoparticle dipole momentum operator
@@
where $p$ is the dipole momentum matrix element, $\hat{b}$ is a Bose-operator describing collective (plasmonic) oscillations of the nanoparticle electrons in the electric field. Later we express $p$ in terms of the polarisability of the nanoparticles [25]. For reasons of symmetry, we assume the same directions for the nanoparticle dipole and the external field polarisation.
The dipole momentum operator of the $i$th absorber resonance transition is
${{\hat{\vec{d}}}_{i}}=d\left[\left({{{\hat{\sigma}}}_{i}}\vec{e}+{{{\hat{\sigma}}}_{\bot i}}{{{\vec{e}}}_{i}}\right){{e}^{-i{{\omega}_{0}}t}}+h.c.\right],$ (2)
where $d$ is the dipole transition matrix element; absorber operators in Dirac notation ${{\hat{\sigma}}_{i}}={{\left|00\right\rangle}_{i}}{{\left\langle 01\right|}_{i}}$, (or ${{{\hat{\sigma}}}_{\bot i}}={{\left|00\right\rangle}_{i}}{{\left\langle 10\right|}_{i}}$) are transition operators from the high $\left|01\right>_{i}$ (or $\left|10\right>_{i}$) to the low $\left|00\right>_{i}$ states of the absorber $i$ parallel to $\vec{e}$, (or $\vec{e}_{i}$): ${{{\hat{\sigma}}}_{i}}\left|01\right>_{i}=\left|00\right>_{i}$ and ${{{\hat{\sigma}}}_{\perp i}}\left|10\right>_{i}=\left|00\right>_{i}$.
The Hamiltonian of the particle and $N_{m}$ absorbers in the electric field $\hat{\vec{E}}$ is
@@
The first term in (3) describes the free oscillations of the dipoles; the second term is for the dipole interaction of the particle and absorbers with the external field; the third term ${{\hat{V}}_{int}}$ describes the interaction of the particle and absorbers through the nanoparticle near field (the LPR field), and the last term in (3) describes the dissipation and the interaction of the particle and absorbers with baths.
We assume that the distance $r$ between the absorbers and the particle centre is much smaller than the LPR wavelength. Thus the particle-absorber interaction ${{\hat{V}}_{int}}$ is a quasi-static dipole-dipole interaction. We take the quasi-static dipole-dipole interaction energy from [45], replace the classical dipoles by operators, and write
@@
where ${{{\vec{e}}}_{ri}}$ is the unit vector along the direction from the nanoparticle centre to the ith absorber, as shown in Fig. 1.
@@
where $\hat{a}$ is the field amplitude Bose operator, ${{E}_{0}}=\eta^{-1}\sqrt{{2\pi\hbar{{\omega}_{0}}}/{V}}$, $V$ is the field quantization volume ($V$ is not used in the final results [46]), $\eta$ is the refractive index of the nanoparticle environment [47]. We assume that the distances between the absorbers and the particle centre are much smaller than the wavelength of the external field, so that the phase (assumed to be zero) of the external field is the same for all absorbers and the particle.
Substituting the expressions (1), (2) into the sum in Eq (4), dropping the fast oscillating terms $\sim e^{\pm 2i\omega_{0}t}$, we get
@@
Expression (6) can be simplified for a large number of absorbers $N_{m}\gg 1$. Summing over $i$, we approximate ${{\cos}^{2}}{{\theta}_{i}}\approx 1/2$ and neglect the product ${{{\hat{\sigma}}}_{\bot i}}\cos{{\theta}_{i}}\sin{{\theta}_{i}}$ repeatedly changing the sign with $i$. After such a simplification of Eq. (6), the Hamiltonian (3) written in the interaction picture [48] and RWA approximation [31] becomes
@@
where ${{\Omega}_{p}}={p{{E}_{0}}}/{\hbar}$, ${{\Omega}_{m}}={d{{E}_{0}}}/{\hbar}$ and ${{\Omega}_{int}}={pd}/2\hbar{{r}^{3}}$. The Hamiltonian (7) corresponds to the approximation when the polarisation of the absorber transitions is parallel to the polarisation of the external field.
2.3 Equations of motion
We write Heisenberg the equations of motion $i\hbar\dot{\hat{A}}=\left[\hat{A},H\right]$ where $\hat{A}$ is an operator from the Hamiltonian (7). The equation for the plasmonic oscillation Bose operator $\hat{b}$ is